Snow has an average density of 15kg/m3. What is the density of the snow in g/cm3?


1
Expert's answer
2021-08-27T08:26:19-0400
"\rm 15\:\frac{kg}{m^3}=15\:\frac{10^3\: g}{10^6\:cm^3}=0.015\:\frac{g}{cm^3}"